3) The mean height of men in the USA (ages 20-29) is 69.9 inches. A random sample of 60 men in this age group is selected. What is the probability that the mean height for the sample is less than 70 inches? 2 3
Answer #1

P(\bar x < 70)

= P((\bar x - \mu )/(\sigma/\sqrt n) < (70 - \mu )/(\sigma/\sqrt n))

= P(Z < (70 - 69.9)/(3/\sqrt 60))

= P(Z < 0.26)

= 0.6026